(1) Let X be exponential random variable with λ = 1.

(a) (4 pts) Define Y = √ X. Specify the support of Y and find its density. Show all of your work and computations.

(b) (6 pts) Define Z = X^2 + 2X. Specify the support of Z and find its density. Show all of your work and computation
Answer #1

200 2 fyly) = f Fyly) = 24042 Ans Let y be an exponential random variable with d=1 Then density of x is be f(x) = { ēm 07 els2 Since Z=4²+2x 3) (x+1) - Ztl 3) X= V2+1 -1 so thal dx ļ (2+1 3 2 dz So the density function of ? is fzlz)= fx(x = 17+1 -1)